OpenStudy (anonymous):

Rewrite with only sin x and cos x. cos 3x

zepdrix (zepdrix):

Hi there! \(\Large\bf \color{#CC0033}{\text{Welcome to OpenStudy! :)}}\)

zepdrix (zepdrix):

\[\Large\rm \cos(3x)\]So we need to somehow write this in terms of an angle x, yes? 3x no beuno.

OpenStudy (anonymous):

I think so, I'm just not sure. Would you like me to give you the answer choices?

zepdrix (zepdrix):

Let's start with our Angle Sum Identity for Cosine,\[\Large\rm \cos(a+b)=\cos(a)\cos(b)-\sin(a)\sin(b)\]\[\Large\rm \cos(3x)=\cos(x+2x)\]Understand how we can apply this rule to our problem?\[\Large\rm \cos(x+2x)=?\]

zepdrix (zepdrix):

Yah the answer choices might be a good idea, sometimes they simplify it in a weird way.

OpenStudy (anonymous):

cos x - 4 cos x sin^2x -sin^3x + 2 sin x cos x -sin^2x + 2 sin x cos x 2 sin^2x cos x - 2 sin x cos x

zepdrix (zepdrix):

Ok good, those will help us when we're getting close to a solution. How bout this first step? Understand how to apply it? :o

OpenStudy (anonymous):

Ummm yeah, so\[\cos(2x+x)=\cos2xcosx+\sin2xsinx\] ????

zepdrix (zepdrix):

\[\Large\rm \cos(x+2x)=\cos(x)\cos(2x)+\sin(x)\sin(2x)\]Ok good. I'm going to keep the brackets up for now.

OpenStudy (anonymous):

Alright that's fine.

zepdrix (zepdrix):

We could apply the formula again if we wanted to,\[\Large\rm \cos(2x)=\cos(x+x)\]But that's going to be a little more tedious. Let's instead apply our Double Angle Formulas.\[\Large\rm \cos(2x)=1-2\sin^2x\]\[\Large\rm \sin(2x)=2\sin(x)\cos(x)\]

zepdrix (zepdrix):

There are like... 3 different ways to write your Cosine Double Angle Formula, I chose this one because all of your answer choices include powers of sinx. So this is probably the form we want. Otherwise it'll leave us with an extra step or two at the end.

OpenStudy (anonymous):

Okay that makes sense.

zepdrix (zepdrix):

\[\Large\rm \cos x\color{royalblue}{\cos(2x)}+\sin x\color{orangered}{\sin(2x)}\]Applying our identities,\[\Large\rm \cos x\color{royalblue}{\left(1-2\sin^2x\right)}+\sin x\color{orangered}{(2\sin x \cos x)}\]Something like that, yes? :o

zepdrix (zepdrix):

You'll need to distribute that cosine to each term in the blue brackets.

zepdrix (zepdrix):

And then you'll need to combine like-terms.

OpenStudy (anonymous):

Do I do the same for the sinx? Distribute sinx to each term in the red brackets?

zepdrix (zepdrix):

Careful... In the orange brackets, notice there is no addition/subtraction sign. So in those brackets, that is a SINGLE TERM. You would distribute the sinx ONE TIME to that block of stuff. Example:\[\Large\rm 2(3\cdot4)\ne2\cdot3\cdot2\cdot4\]\[\Large\rm 2(3\cdot4)=2\cdot3\cdot4\]We distribute one 2 in this example^

OpenStudy (anonymous):

Alrighty so then\[cosx-2\sin^2xcosx+2sinx^2xcosx\]

zepdrix (zepdrix):

Ahhh we made a mistake earlier. Our Double Angle Identity for Cosine should have given us subtraction here:\[\Large\rm \cos(x+2x)=\cos(x)\cos(2x)\color{red}{+}\sin(x)\sin(2x)\]

zepdrix (zepdrix):

So we actually have:\[\Large\rm \cos x\color{royalblue}{\left(1-2\sin^2x\right)}-\sin x\color{orangered}{(2\sin x \cos x)}\]And then yes, distributing would give us:\[\Large\rm \cos x-2\sin^2x \cos x-2\sin^2x \cos x\]

OpenStudy (anonymous):

Ahhh, okay! So are there a few more steps until we get to the answer???

zepdrix (zepdrix):

One more step it looks like :) Hmm we can combine like-terms from here.

OpenStudy (anonymous):

OH, so the answer is the first answer choice. \[cosx-4cosxsin^2x\]

zepdrix (zepdrix):

Yayyy good job \c:/

OpenStudy (anonymous):

Thank you for all your help :))

zepdrix (zepdrix):

np